# Disentangling the Halo: Joint Model for Measurements of the Kinetic Sunyaev-Zeldovich Effect and Galaxy-Galaxy Lensing

### v1

@@ 5 Conclusions
[…]
@@
Going forward, there are many avenues to build upon this analysis. First, new measurements of the stacked kSZ profile from ACT-DR6 and DESI-Y1 data using photometric (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2024; Hadzhiyska et al., 2024b) and spectroscopic (Ried Guachalla et al., 2025a) redshifts have significantly higher SNR than the measurements from Schaan et al. (2021) (analysed in this work) with the ability to measure the signal in a more narrow range of redshift and mass. In the near future, Simons Observatory (SO) will bring better CMB measurements to accompany the DESI spectroscopic survey and yield better kSZ profile measurements. At the same time, upcoming WL surveys such as Euclid, LSST, and Roman will provide substantially more powerful WL measurements. Additionally, when combined, the kSZ effect and the tSZ effect can probe the thermodynamic profiles around galaxies (Amodeo et al., 2021), when combined with additional information from GGL one can have unprecedented constraining power for testing baryonic feedback models. Together, these measurements will enable a better understanding of baryonic feedback, halo masses, and cosmology.
@@ Acknowledgements
The stacked kSZ profile measurements used in this work can be made available upon request and are from the original work of Schaan et al. (2021). The GGL measurements used in this analysis are publicly available on this github repository and were originally made by Amon et al. (2023). [^7] https://github.com/aamon/Galaxy-galaxy-lensing-KiDS-DES-HSC Our modelling code glasz is built upon the pyccl Chisari et al. (2019b) and Mop-c-GT Amodeo et al. (2021) packages and is publicly available at the glasz homepage. [^8] https://github.com/James11222/glasz We request that any uses of the glasz package cite the DOI listed on the package homepage.

### v2

@@ 5 Conclusions
    the kSZ effect serves as a complementary probe to GGL. By doing a joint fit to kSZ+GGL data, we can alleviate degeneracies in the model that exist when fitting to a single dataset (kSZ or GGL) as seen in Figures 2 and 3. When our fiducial model is fit to both kSZ and GGL measurements, the baryon density profile is tightly constrained at all relevant scales $r\approx 0.3-50$ Mpc/$h$ (in comoving units) as seen in the left panel of Figure 4. By having a tight constraint on the baryon density profile around the outskirts of the galactic halo, we determine empirically how baryons suppress $\Delta\Sigma$ at small scales as seen in the right panel of Figure 4. This work opens an avenue for using small scales in future cosmological weak lensing analyses. Additionally, not including baryons in a model of GGL can shift halo mass estimates by $\sim 20\%$ compared to a model that includes baryons fit to kSZ+GGL measurements as seen in Figure 5.
@@
Going forward, there are many avenues to build upon this analysis. First, new measurements of the stacked kSZ profile from ACT-DR6 and DESI-Y1 data using photometric (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2024; Hadzhiyska et al., 2024b) and spectroscopic (Ried Guachalla et al., 2025b) redshifts have significantly higher SNR than the measurements from Schaan et al. (2021) (analysed in this work) with the ability to measure the signal in a more narrow range of redshift and mass. In the near future, Simons Observatory (SO) will bring better CMB measurements to accompany the DESI spectroscopic survey and yield better kSZ profile measurements. At the same time, upcoming WL surveys such as Euclid, LSST, and Roman will provide substantially more powerful WL measurements. Additionally, when combined, the kSZ effect and the tSZ effect can probe the thermodynamic profiles around galaxies (Amodeo et al., 2021), when combined with additional information from GGL one can have unprecedented constraining power for testing baryonic feedback models. Together, these measurements will enable a better understanding of baryonic feedback, halo masses, and cosmology.
@@ Acknowledgements
The stacked kSZ profile measurements used in this work can be made available upon request and are from the original work of Schaan et al. (2021). The GGL measurements used in this analysis are publicly available on this github repository and were originally made by Amon et al. (2023). [^7] https://github.com/aamon/Galaxy-galaxy-lensing-KiDS-DES-HSC Our modelling code glasz is built upon the pyccl Chisari et al. (2019a) and Mop-c-GT Amodeo et al. (2021) packages and is publicly available at the glasz homepage. [^8] https://github.com/James11222/glasz We request that any uses of the glasz package cite the DOI listed on the package homepage.
@@ Appendix B Differences from Amodeo Halo Model
Appendix C Degeneracy in Baryon Profile Parameters
The 2 baryon density profile parameters $x_{\rm c}$ and $\beta$ are highly degenerate with each other. While we have defined the core scale fraction $x_{\rm c}$ to span $[0,1]$ based off of the definition proposed previously in precursory work (Battaglia, 2016; Amodeo et al., 2021), this is simply a choice in the model. In Smith et al. (2018) the core scale fraction is fixed to $x_{\rm c}=2$ rather than 0.5. A higher value of $x_{\rm c}$ implies that the turnover in the density profile occurs further out into the halo which requires a higher (steeper) value $\beta$ for the baryon density to remain equivalent in the outskirts of the halo.
To better understand this degeneracy between $x_{\rm c}$ and $\beta$, we fit the data with expanded priors on the $x_{\rm c}$ parameter. We find that only one of the two parameters can be constrained at a time. More specifically, when the MCMC chains reach a prior wall for one of the two parameters, the other parameter will be constrained. If the upper bound on the prior for $x_{\rm c}$ is 5, then $\beta$ will remain unconstrained. If the upper bound on the prior for $\beta$ is expanded to 10 in this case, then $\beta$ will be constrained while $x_{\rm c}$ remains unconstrained (the posteriors look similar to Figure 3 only shifted to higher values). Fixing $x_{\rm c}=2.0$ necessarily shifts the marginalized constraint on the outer power law slope to $\beta=3.60\pm 0.19$ while leaving $A_{\rm 2h}$ and $M_{\rm halo}$ unchanged. Ultimately, in order to make a fair comparison to the simulations of Battaglia (2016), we choose to keep the fiducial prior $\mathcal{U}(0.0,1.0)$ for $x_{\rm c}$ which is consistent with the model definition.